OpenStudy (anonymous):

I got 2006.25 on this question but I don't think the answer is correct

OpenStudy (anonymous):

Here's the question

OpenStudy (anonymous):

what numbers did you use?

OpenStudy (anonymous):

I did 306/321*100=95.3% Second, I wanted to find the decrease of the percent 321 which is 16 so 321-306=16 (321/16)*100=2006.25 or 8025/4

OpenStudy (anonymous):

you did your steps backwards. Find the decrease first then divide by the first value. \[\frac{ 321-306 }{ 321 }\times100\]

OpenStudy (anonymous):

4.672

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The answer was 5.8